OpenStudy (anonymous):

Use the method of sustiution to solve the system of linear equations. 6a=6-b 6a=3-b

hartnn (hartnn):

u have value of 6a from 1st equation just substitute in 2nd equation like this 6-b=3-b now can u solve for b?

OpenStudy (anonymous):

it would be 3b

hartnn (hartnn):

are u sure about your question??please verify...

OpenStudy (anonymous):

would it be 9b because you multiply

OpenStudy (anonymous):

no... are you sure thats the question?

OpenStudy (anonymous):

I have to find out if there are any solutions no solutions and if there is what is the ordered pair

OpenStudy (anonymous):

then there are no solutions

OpenStudy (anonymous):

yup that is the question us the method of substiution to solve the system of linear equations

OpenStudy (anonymous):

there are no solutions

OpenStudy (anonymous):

your question is not correct for any value of a and b.

OpenStudy (anonymous):

Okay thank you, that is what I thought but I got a little confused on it

hartnn (hartnn):

then as @Mayb40 said there will be no solution because for no value b, 6-b=3-b satisfies...